the length of a rectangular garden is 5 m greater than the width. the area is 104m^2. find the dimensions of the garden.
Answer by jim_thompson5910(35256)

A = LW

104 = (W+5)W

104 = W(W+5)

W^2 + 5W - 104 = 0

Now use the quadratic formula to solve

W = (-b+-sqrt(b^2-4ac))/(2a)

W = (-(5)+-sqrt((5)^2-4(1)(-104)))/(2(1))

W = (-5+-sqrt(25-(-416)))/(2)

W = (-5+-sqrt(441))/2

W = (-5+sqrt(441))/2 or W = (-5-sqrt(441))/2

W = (-5+21)/2 or W = (-5-21)/2

W = 16/2 or W = -26/2

W = 8 or W = -13

Throw out the negative solution to be left with the only solution of W = 8

So the width is 8 m and length is 13 m (since L = W+5 = 8+5 = 13)